# State Estimation Using Sparse DEIM and Recurrent Neural Networks

### v2

@@ Abstract
Sparse Discrete Empirical Interpolation Method (S-DEIM) was recently proposed for state estimation in dynamical systems when only a sparse subset of the state variables can be observed. The S-DEIM estimate involves a kernel vector whose optimal value is inferred through a data assimilation algorithm. This data assimilation step suffers from two drawbacks: (i) It requires the knowledge of the governing equations of the dynamical system, and (ii) It is not generally guaranteed to converge to the optimal kernel vector. To address these issues, here we introduce an equation-free S-DEIM framework that estimates the optimal kernel vector from sparse observational time series using recurrent neural networks (RNNs). We show that the recurrent architecture is necessary since the kernel vector cannot be estimated from instantaneous observations. But RNNs, which incorporate the past history of the observations in the learning process, lead to nearly optimal estimations. We demonstrate the efficacy of our method on three numerical examples with increasing degree of spatiotemporal complexity: a conceptual model of atmospheric flow known as the Lorenz-96 system, the Kuramoto–Sivashinsky equation, and the Rayleigh–Bénard convection. In each case, the resulting S-DEIM estimates are satisfactory even when a relatively simple RNN architecture, namely the reservoir computing network, is used.
@@ 1 Introduction
where $\mathbf{u}(t)$ is the state to be estimated, $\mathbf{y}(t)$ represents the observations at time $t$, $\Phi$ and $\Theta$ are appropriate matrices as defined in Section 2, and the superscript $+$ denotes the pseudo inverse of a matrix. The first term in (1) can immediately be computed from the observations $\mathbf{y}(t)$. The second term involves the kernel vector $\mathbf{z}(t)$ which belongs to the null space of $\Theta$ and whose optimal value needs to be estimated. The (Q)-DEIM estimate uses to the trivial kernel vector $\mathbf{z}(t)=\mathbf{0}$. In contrast, the S-DEIM estimate uses a nonzero optimal kernel vector.
@@
To address these issues, here we develop an equation-free method to approximate the optimal kernel vector in S-DEIM. Our method relies on a recurrent neural network to learn the optimal kernel vector $\mathbf{z}(t)$ from time series of the observational data $\mathbf{y}(s)$ with $s\leq t$, leveraging the information encoded in the past history of the observations. We demonstrate, with several numerical examples, that the resulting state estimations are nearly optimal.
@@ 1.1 Related work
State estimation: As the name suggests, DEIM can be used for interpolation purposes. In the context of dynamical systems, this interpolation amounts to state estimation from incomplete observations [38, 33, 7]. Let the state variable be $N$-dimensional and assume that only an $r$-dimensional subset of them can be observed. As we review in Section 2, DEIM can be used to approximate the remaining $N-r$ variables within a reduced subspace spanned by $m$ modes.
@@
Sensor placement: The quality of DEIM-based state estimation depends significantly on the location of the sensors which gather the observational data. For instance, if the sensor locations are assigned at random, the DEIM estimation error is typically very large. Drmač and Gugercin [12] proposed the column-pivoted QR (CPQR) factorization as a greedy algorithm for approximating the set of optimal sensor locations. The resulting method is dubbed Q-DEIM and is reviewed in Section 2.2 below. In the overdetermined case where $r>m$, CPQR only returns the optimal locations for the first $m$ sensors. Hendryx et al. [19] proposed extended DEIM (E-DEIM) to approximate the optimal location of the remaining $r-m$ sensors; also see [28].
@@ 2 Sparse DEIM
We consider a dynamical system whose state at time $t$ is denoted by $\mathbf{u}(t)\in\mathbb{R}^{N}$. The state can, for instance, denote the solution of an ordinary differential equation (ODE) or a spatially discretized solution of a time-dependent partial differential equation (PDE).
@@
corresponding to the coefficients $\hat{\boldsymbol{\alpha}}=\Phi_{m}^{\top}\mathbf{u}$ which minimize the truncation error $\mathcal{E}_{m}(\boldsymbol{\alpha}):=|\Phi_{m}\boldsymbol{\alpha}-\mathbf{u}|$, where ||$\cdot$|| ⋅ | denotes the standard Euclidean norm of a vector. The truncation error $\mathcal{E}_{m}$ is a monotonically decreasing function of $m$, i.e., using more modes leads to better approximations. Unfortunately, however, the best fit approximation (4) is not computable since it relies on the full state $\mathbf{u}$.
To overcome this issue, DEIM determines the coefficients $\boldsymbol{\alpha}$ by minimizing the discrepancy between the observations $\mathbf{y}$ and the reconstructed state at the sensor locations. To this end, we define the observation error,
$\mathcal{E}_{obs}(\boldsymbol{\alpha})=|S_{r}^{\top}\Phi_{m}\boldsymbol{\alpha}-\mathbf{y}|.$ (5)
@@
for an arbitrary $\mathbf{z}\in\mathcal{N}[S_{r}^{\top}\Phi_{m}]$, where the superscript $+$ denotes the Moore-Penrose pseudo-inverse and $\mathcal{N}$ denotes the null space of a linear operator. The matrix $\Theta$ in equation (1) is equal to $S_{r}^{\top}\Phi_{m}$.
@@
Figure 1: State space geometry of DEIM estimation with $r<m$. The truncated basis $\{\boldsymbol{\phi}_{1},\cdots,\boldsymbol{\phi}_{m}\}$ spans the subspace $\mathcal{R}[\Phi_{m}]\subset\mathbb{R}^{N}$. The variables denote the true state $\mathbf{u}$, the best fit $\hat{\mathbf{u}}$, DEIM estimate $\widetilde{\mathbf{u}}(\mathbf{0})$, and the optimal S-DEIM estimate $\widetilde{\mathbf{u}}(\hat{\mathbf{z}})$. The subspace $\mathcal{R}[\Phi_{m}]$ admits the orthogonal decomposition $\mathcal{R}[\Phi_{m}]=\mathcal{R}[(S_{r}^{\top}\Phi_{m})^{+}]\oplus\Phi\mathcal{N}[S_{r}^{\top}\Phi_{m}]$.
@@
$\mathcal{E}_{tot}(\mathbf{z})=|\widetilde{\mathbf{u}}(\mathbf{z})-\mathbf{u}|.$ (8)
@@
However, this optimal kernel vector is not computable since it requires the knowledge of the full state $\mathbf{u}$. In Section 3, we introduce a machine learning method for estimating the optimal kernel vector from observational time series. But first we briefly review the common computational methods for determining the basis matrix $\Phi_{m}$ (Section 2.1) and the selection matrix $S_{r}$ (Section 2.2).
@@ 2.2 Sensor placement: Column-pivoted QR decomposition
Drmač and Gugercin [12] proposed the column-pivoted QR (CPQR) decomposition for sensor placement in DEIM. The resulting method is referred to as Q-DEIM. CPQR was first proposed for the special case $r=m$ in [12] and later extended to the case with $r<m$ in [24]. In the present work, we also use the CPQR algorithm for sensor placement, and therefore review the method briefly in this section.
@@
$\frac{|\widetilde{\mathbf{u}}(\mathbf{0})-\mathbf{u}|}{|\mathbf{u}|}\leq\|(S_{r}^{\top}\Phi_{m})^{+}\|_{2},$ (12)
where $\|\cdot\|_{2}$ denotes the spectral norm (i.e., 2-norm) of a matrix. CPQR seeks to determine the selection matrix $S_{r}$ that minimizes the upper bound $\|(S_{r}^{\top}\Phi_{m})^{+}\|_{2}$.
@@
On the other hand, consider the rank-revealing QR decomposition of $\Phi_{m}^{\top}\in\mathbb{R}^{m\times N}$ with column pivoting,
@@
In practice, when the selection matrix $S_{r}$ is determined by CPQR, the resulting matrix $S_{r}^{\top}\Phi_{m}$ often has full row rank [16]. In the following, we assume this full-rank property.
Assumption 1.
Let $r<m$. Given a basis matrix $\Phi_{m}$ and the corresponding selection matrix $S_{r}$ obtained from the CPQR decomposition (14), we assume that the matrix $S_{r}^{\top}\Phi_{m}$ has full row rank.
Recall that the pseudo-inverse of a full row rank matrix is a right inverse, so that $(S_{r}^{\top}\Phi_{m})(S_{r}^{\top}\Phi_{m})^{+}=I_{r}$. More specifically, if $A$ has full row rank, then $A^{+}=A^{\top}(AA^{\top})^{-1}$.
@@ 3 S-DEIM Using Recurrent Neural Networks
Under Assumption 1, S-DEIM estimation (7) is an interpolation in the following sense. For any kernel vector $\mathbf{z}$, we have
@@ 3.1 Training the reservoir computing network
Before describing the RC architecture, we first outline the data required for its training. We assume access to a set of high-fidelity data $\mathbf{u}(t_{i})$ at discrete time instances $t_{0}<t_{1}<t_{2}<\cdots<t_{M}$. We further assume that the data is equispaced in time so that $t_{i+1}-t_{i}=\Delta t$. From this high-fidelity data set we compute the corresponding observations $\mathbf{y}(t_{i})=S_{r}^{\top}\mathbf{u}(t_{i})$ and the optimal kernel vectors $\hat{\mathbf{z}}(t_{i})=ZZ^{\top}\Phi_{m}^{\top}\mathbf{u}(t_{i})$. Our aim is to train a neural network whose inputs are the observational time series $\mathbf{y}(t_{i})$ and its outputs are the optimal kernel vectors $\hat{\mathbf{z}}(t_{i})$. However, it is more prudent to train the neural networks with outputs being $\hat{\boldsymbol{\xi}}(t_{i})=Z^{\top}\hat{\mathbf{z}}(t_{i})=Z^{\top}\Phi_{m}^{\top}\mathbf{u}(t_{i})$. Recall that the columns of the kernel matrix $Z$ form an orthonormal basis for $\mathcal{N}[S_{r}^{\top}\Phi_{m}]$, and therefore the vector $\hat{\boldsymbol{\xi}}(t_{i})\in\mathbb{R}^{m-r}$ denotes the coordinates of $\hat{\mathbf{z}}(t_{i})\in\mathbb{R}^{m}$ in that basis. Since there is a one-to-one correspondence between $\hat{\mathbf{z}}(t_{i})$ and $\hat{\boldsymbol{\xi}}(t_{i})$, replacing the network output with $\hat{\boldsymbol{\xi}}(t_{i})$ does not lead to any loss of information; the optimal kernel vector can be computed uniquely as $\hat{\mathbf{z}}(t_{i})=Z\hat{\boldsymbol{\xi}}(t_{i})$. Yet, the vector $\hat{\boldsymbol{\xi}}(t_{i})$ has dimension $m-r$, which is smaller than the dimension of the optimal kernel vector $\hat{\mathbf{z}}(t_{i})\in\mathbb{R}^{m}$. This lower dimensionality leads to a faster and more accurate training process.
@@
What sets RC networks apart from most other neural networks is the fact that its input weights $W_{in}$, its reservoir weights $W_{R}$ and the biases $\mathbf{b}$ of the network are prescribed at random, as opposed to being determined through an optimization process. The only trained part of the network is the output weights $W_{out}$. As we will see shortly, there exists a simple closed form expression for the optimal output weights. As such, training RC networks is significantly faster than other recurrent neural networks such as long short-term memory (LSTM) networks.
@@
The updates (18) are designed so that the reservoir state $\mathbf{r}(t_{i})$ synchronizes with the input time series of the observations $\mathbf{y}(t_{i})$. As a result, the long-term dynamics of the reservoir should be independent of the initial state $\mathbf{r}(t_{0})$ of the reservoir. This feature is referred to as the echo state property [23, 32]. To ensure that the echo state property is satisfied, we use a commonly used heuristic, namely that the spectral radius of the reservoir matrix $W_{R}$ is less than one [47]. We enforce this property by first defining a sparse matrix $W_{R}$ whose nonzero entries are randomly drawn from the uniform distribution $U([-0.5,0.5])$. We subsequently rescale this matrix to ensure the desired spectral radius. Similarly, the entries of the input matrix $W_{in}$ and the biases $\mathbf{b}$ are drawn at random from the uniform distribution $U([-0.5,0.5])$. The spectral radius of the input weights $W_{in}$ does not play an important role, and therefore no rescaling is needed for $W_{in}$.
To summarize, the observational time series $\mathbf{y}(t_{i})$ act as an external forcing term that drives the reservoir dynamics (18). The reservoir state $\mathbf{r}(t_{i})$ is mapped to the outputs through the weight matrix $W_{out}$. The outputs are then mapped to the kernel vector $\mathbf{z}(t_{i})$ through the kernel matrix $Z$. Finally, we use this kernel vector in (7) to arrive at the S-DEIM estimate $\widetilde{\mathbf{u}}(\mathbf{z}(t_{i}))$. Across the examples presented in Section 4, the training process takes only a few seconds. Once the RC network is trained, the S-DEIM estimate is obtained within milliseconds.

### v3

@@ Abstract
Sparse Discrete Empirical Interpolation Method (S-DEIM) was recently proposed for state estimation in dynamical systems when only a sparse subset of the state variables can be observed. The S-DEIM estimate involves a kernel vector whose optimal value is inferred through a data assimilation algorithm. This data assimilation step suffers from two drawbacks: (i) It requires the knowledge of the governing equations of the dynamical system, and (ii) It is not generally guaranteed to converge to the optimal kernel vector. To address these issues, here we introduce an equation-free S-DEIM framework that estimates the optimal kernel vector from sparse observational time series using recurrent neural networks (RNNs). We show that the recurrent architecture is necessary since the kernel vector cannot be estimated from instantaneous observations. But RNNs, which incorporate the past history of the observations in the learning process, lead to nearly optimal estimations. We demonstrate the efficacy of our method on three numerical examples with increasing degree of spatiotemporal complexity: a conceptual model of atmospheric flow known as the Lorenz-96 system, the Kuramoto–Sivashinsky equation, and the Rayleigh–Bénard convection. In each case, the resulting S-DEIM estimates are satisfactory even when a relatively simple RNN architecture, namely the reservoir computing network, is used. More specifically, our RNN-based S-DEIM state estimations reduce the relative error between 42% and 58% when compared to Q-DEIM which ignores the kernel vector by setting it equal to zero.
@@ 1 Introduction
where $\mathbf{u}(t)$ is the state to be estimated, $\mathbf{y}(t)$ represents the observations at time $t$, $\Phi$ and $\Theta$ are appropriate matrices as defined in Section 2, and the superscript ++ denotes the pseudo inverse of a matrix. The first term in (1) can immediately be computed from the observations $\mathbf{y}(t)$. The second term involves the kernel vector $\mathbf{z}(t)$ which belongs to the null space of $\Theta$ and whose optimal value needs to be estimated. The (Q)-DEIM estimate uses to the trivial kernel vector $\mathbf{z}(t)=\mathbf{0}$. In contrast, the S-DEIM estimate uses a nonzero optimal kernel vector.
@@
To address these issues, here we develop an equation-free method to approximate the optimal kernel vector in S-DEIM. Our method relies on a Recurrent Neural Network (RNN) to learn the optimal kernel vector $\mathbf{z}(t)$ from time series of the observational data $\mathbf{y}(s)$ with $s\leq t$, leveraging the information encoded in the past history of the observations. As such, the main innovation in this paper is the use of RNNs to estimate the optimal kernel vector in S-DEIM, without relying on the governing equations of the dynamical systems. We demonstrate, with several numerical examples, that the resulting state estimations are nearly optimal.
@@ 1.1 Related work
State estimation: As the name suggests, DEIM can be used for interpolation purposes. In the context of dynamical systems, this interpolation amounts to state estimation from incomplete observations [22, 6, 4]. Let the state variable be $N$-dimensional and assume that only an $r$-dimensional subset of them can be observed. As we review in Section 2, DEIM can be used to approximate the remaining NrN-r variables within a reduced subspace spanned by $m$ modes.
@@
Sensor placement: The quality of DEIM-based state estimation depends significantly on the location of the sensors which gather the observational data. For instance, if the sensor locations are assigned at random, the DEIM estimation error is typically very large. Drmač and Gugercin [7] proposed the column-pivoted QR (CPQR) factorization as a greedy algorithm for approximating the set of optimal sensor locations. The resulting method is dubbed Q-DEIM and is reviewed in Section 2.3 below. In the overdetermined case where $r>m$, CPQR only returns the optimal locations for the first $m$ sensors. Hendryx et al. [28] proposed extended DEIM (E-DEIM) to approximate the optimal location of the remaining rmr-m sensors; also see [29].
@@ 2 Sparse DEIM
We consider a dynamical system whose state at time $t$ is denoted by $\mathbf{u}(t)\in\mathbb{R}^{N}$. The state can, for instance, denote the solution of an ordinary differential equation (ODE) or a spatially discretized solution of a time-dependent partial differential equation (PDE). Table 1 summarizes the key terms and notation used in this paper.
Table 1: Glossary of key terms and notation.
Symbol Description
NN State Space Dimension
$r$ Number of Sensors
$m$ Number of Modes (Basis Functions)
$\mathbf{u}(t)\in\mathbb{R}^{N}$ Unknown State
$\mathbf{y}(t)\in\mathbb{R}^{r}$ Known Observations
$S_{r}\in\mathbb{R}^{N\times r}$ Selection Matrix, $\mathbf{y}(t)=S_{r}^{\top}\mathbf{u}(t)$
$\Phi_{m}=[\boldsymbol{\phi}_{1}|\cdots|\boldsymbol{\phi}_{m}]\in\mathbb{R}^{N\times m}$ Basis Matrix
$\widetilde{\mathbf{u}}(\mathbf{z}(t))\in\mathbb{R}^{N}$ S-DEIM Estimate
$\hat{\mathbf{u}}(t)\in\mathbb{R}^{N}$ Best Fit, $\hat{\mathbf{u}}(t)=\Phi_{m}\Phi_{m}^{\top}\mathbf{u}(t)$
$\hat{\mathbf{z}}(t)\in\mathbb{R}^{m}$ Optimal Kernel Vector, $\hat{\mathbf{z}}(t)=ZZ^{\top}\Phi_{m}^{\top}\mathbf{u}(t)$
$\mathcal{E}_{m}(\boldsymbol{\alpha}):=\|\Phi_{m}\boldsymbol{\alpha}-\mathbf{u}\|$ Truncation Error
$\mathcal{E}_{obs}(\boldsymbol{\alpha})=\|S_{r}^{\top}\Phi_{m}\boldsymbol{\alpha}-\mathbf{y}\|$ Observation Residual
@@
corresponding to the coefficients $\hat{\boldsymbol{\alpha}}=\Phi_{m}^{\top}\mathbf{u}$ which minimize the truncation error $\mathcal{E}_{m}(\boldsymbol{\alpha}):=\|\Phi_{m}\boldsymbol{\alpha}-\mathbf{u}\|$, where $\|\cdot\|$ denotes the standard Euclidean norm of a vector. The truncation error $\mathcal{E}_{m}$ is a monotonically decreasing function of $m$, i.e., using more modes leads to better approximations. Unfortunately, however, the best fit approximation (4) is not computable since it relies on the full state $\mathbf{u}$.
To overcome this issue, DEIM determines the coefficients $\boldsymbol{\alpha}$ by minimizing the discrepancy between the observations $\mathbf{y}$ and the reconstructed state at the sensor locations. To this end, we define the observation residual,
$\mathcal{E}_{obs}(\boldsymbol{\alpha})=\|S_{r}^{\top}\Phi_{m}\boldsymbol{\alpha}-\mathbf{y}\|.$ (5)
@@
for an arbitrary $\mathbf{z}\in\mathcal{N}[S_{r}^{\top}\Phi_{m}]$, where the superscript ++ denotes the Moore-Penrose pseudo-inverse and $\mathcal{N}$ denotes the null space of a linear operator. The matrix $\Theta$ in equation (1) is equal to $S_{r}^{\top}\Phi_{m}$.
@@
[…]
@@
In this underdetermined case, (Q)-DEIM estimate, which uses the trivial kernel vector $\mathbf{z}=\mathbf{0}$, is quite restrictive. As illustrated in figure 1, the DEIM estimate $\widetilde{\mathbf{u}}(\mathbf{0})$ belongs to the range of $(S_{r}^{\top}\Phi_{m})^{+}$, which is a lower-dimensional subspace of $\mathcal{R}[\Phi_{m}]$. As such, the DEIM estimate is constrained to $\mathcal{R}[(S_{r}^{\top}\Phi_{m})^{+}]$ and cannot take advantage of the full expressivity of the basis $\Phi_{m}$. On the other hand, the S-DEIM estimates, which use a non-zero $\mathbf{z}$, are free of this constraint and can explore the full range of the basis matrix $\Phi_{m}$ more freely.
@@
$\mathcal{E}_{tot}(\mathbf{z})=\|\widetilde{\mathbf{u}}(\mathbf{z})-\mathbf{u}\|.$ (8)
@@
However, this optimal kernel vector is not directly computable from observations; it requires the knowledge of the full state $\mathbf{u}$. In Section 3, we introduce a machine learning method for estimating the optimal kernel vector from observational time series. But first we briefly review the S-DEIM error estimate (Section 2.1) as well as the common computational methods for determining the basis matrix $\Phi_{m}$ (Section 2.2) and the selection matrix $S_{r}$ (Section 2.3).
2.1 S-DEIM error
Let $\widetilde{\mathbf{u}}(\mathbf{z})$ denote the S-DEIM estimate (7) with an arbitrary kernel vector $\mathbf{z}\in\mathcal{N}[S_{r}^{\top}\Phi_{m}]$. We would like to quantify the estimation error $\|\widetilde{\mathbf{u}}(\mathbf{z})-\mathbf{u}\|$. It was shown in [8, Theorem 1] that this error can be decomposed into three terms,
$\|\widetilde{\mathbf{u}}(\mathbf{z})-\mathbf{u}\|^{2}=\|\mathbf{u}-\hat{\mathbf{u}}\|^{2}+\|\mathbb{D}(\mathbf{u}-\hat{\mathbf{u}})\|^{2}+\|\mathbf{z}-\hat{\mathbf{z}}\|^{2},\quad\forall\mathbf{z}\in\mathcal{N}[S_{r}^{\top}\Phi_{m}],$ (11)
where $\mathbb{D}=\Phi_{m}(S_{r}^{\top}\Phi_{m})^{+}S_{r}^{\top}$, $\hat{\mathbf{u}}=\Phi_{m}\Phi_{m}^{\top}\mathbf{u}$ is the best fit, and $\hat{\mathbf{z}}=ZZ^{\top}\Phi_{m}^{\top}\mathbf{u}$ is the optimal kernel vector (10). We refer to [8] for the derivation of this identity.
The first term on the right-hand side of (11) involves the truncation error $\|\mathbf{u}-\hat{\mathbf{u}}\|$, reflecting the fact that the basis matrix $\Phi_{m}$ contains a truncated basis of $\mathbb{R}^{N}$. This error can be decreased by increasing the number of modes $m$.
[…]
Finally, the most relevant to our discussion is the third term $\|\mathbf{z}-\hat{\mathbf{z}}\|$ involving the kernel vector $\mathbf{z}$. Our main goal in this paper is to estimate and use the optimal kernel vector $\hat{\mathbf{z}}$ in order to minimize this component of the error as much as possible. (Q)-DEIM, which uses the trivial kernel $\mathbf{z}=\mathbf{0}$, is bound to at least incur an error of $\|\hat{\mathbf{z}}\|$, no matter how many modes $m$ are used or how cleverly the sensors are placed.
@@ 2.3 Sensor placement: Column-pivoted QR decomposition
Drmač and Gugercin [7] used the column-pivoted QR (CPQR) decomposition for sensor placement in DEIM. The resulting method is referred to as Q-DEIM. CPQR was first proposed for the special case $r=m$ in [7] and later extended to the case with $r<m$ in [27]. In the present work, we also use the CPQR algorithm for sensor placement, and therefore review the method briefly in this section.
@@
$\frac{\|\widetilde{\mathbf{u}}(\mathbf{0})-\mathbf{u}\|}{\|\mathbf{u}\|}\leq\|(S_{r}^{\top}\Phi_{m})^{+}\|_{2},$ (13)
where $\|\cdot\|_{2}$ denotes the spectral norm (i.e., 2-norm) of a matrix. In the context of DEIM, sensor placement methods seek to determine the selection matrix $S_{r}$ that minimizes the upper bound $\|(S_{r}^{\top}\Phi_{m})^{+}\|_{2}$.
@@
On the other hand, consider a rank-revealing QR decomposition of $\Phi_{m}^{\top}\in\mathbb{R}^{m\times N}$ with column pivoting,
@@
When the selection matrix $S_{r}$ is determined by CPQR, the resulting matrix $S_{r}^{\top}\Phi_{m}$ has full row rank [38]. Recall that the pseudo-inverse of a full row rank matrix is a right inverse, so that $(S_{r}^{\top}\Phi_{m})(S_{r}^{\top}\Phi_{m})^{+}=I_{r}$. More specifically, if $A$ has full row rank, then $A^{+}=A^{\top}(AA^{\top})^{-1}$.
We emphasize that CPQR sensor placement is derived based on the assumption that $\mathbf{z}=\mathbf{0}$; see equation (13). As such, it is not guaranteed to return optimal (or even nearly optimal) sensor locations for S-DEIM state estimation which uses $\mathbf{z}\neq\mathbf{0}$. Nonetheless, in the absence of a tailor-made sensor placement method for S-DEIM, we use CPQR here.
@@ 3 S-DEIM Using Recurrent Neural Networks
When $S_{r}^{\top}\Phi_{m}$ has full row rank, S-DEIM estimation (7) is an interpolation in the following sense. For any kernel vector $\mathbf{z}$, we have the interpolation property [8],
@@
The RC architecture is chosen here solely for its simplicity and low training cost. More complex networks, such as Long Short-Term Memory (LSTM) networks, can certainly be used instead and will likely improve the resulting S-DEIM estimates. Nonetheless, as we show with our numerical results (Section 4), RC-based S-DEIM already improves the state estimation accuracy by 42-58%.
@@ 3.1 Training the reservoir computing network
Before describing the RC architecture, we first outline the data required for its training. We assume access to a set of high-fidelity data $\mathbf{u}(t_{i})$ at discrete time instances $t_{0}<t_{1}<t_{2}<\cdots<t_{M}$. We further assume that the data is equispaced in time so that $t_{i+1}-t_{i}=\Delta t$. From this high-fidelity data set we compute the corresponding observations $\mathbf{y}(t_{i})=S_{r}^{\top}\mathbf{u}(t_{i})$ and the optimal kernel vectors $\hat{\mathbf{z}}(t_{i})=ZZ^{\top}\Phi_{m}^{\top}\mathbf{u}(t_{i})$. Our aim is to train a neural network whose inputs are the observational time series $\mathbf{y}(t_{i})$ and its outputs are the optimal kernel vectors $\hat{\mathbf{z}}(t_{i})$. However, it is more prudent to train the neural networks with outputs being $\hat{\boldsymbol{\xi}}(t_{i})=Z^{\top}\hat{\mathbf{z}}(t_{i})=Z^{\top}\Phi_{m}^{\top}\mathbf{u}(t_{i})$. Recall that the columns of the kernel matrix $Z$ form an orthonormal basis for $\mathcal{N}[S_{r}^{\top}\Phi_{m}]$, and therefore the vector $\hat{\boldsymbol{\xi}}(t_{i})\in\mathbb{R}^{m-r}$ denotes the coordinates of $\hat{\mathbf{z}}(t_{i})\in\mathbb{R}^{m}$ in that basis. Since there is a one-to-one correspondence between $\hat{\mathbf{z}}(t_{i})$ and $\hat{\boldsymbol{\xi}}(t_{i})$, replacing the network output with $\hat{\boldsymbol{\xi}}(t_{i})$ does not lead to any loss of information; the optimal kernel vector can be computed uniquely as $\hat{\mathbf{z}}(t_{i})=Z\hat{\boldsymbol{\xi}}(t_{i})$. Yet, the vector $\hat{\boldsymbol{\xi}}(t_{i})$ has dimension mrm-r, which is smaller than the dimension of the optimal kernel vector $\hat{\mathbf{z}}(t_{i})\in\mathbb{R}^{m}$. This lower dimensionality leads to a faster and more accurate training process.
@@
What sets RC networks apart from most other neural networks is the fact that its input weights $W_{in}$, its reservoir weights $W_{R}$ and the biases $\mathbf{b}$ of the network are prescribed at random, as opposed to being determined through an optimization process. The only trained part of the network is the output weights $W_{out}$. As we will see shortly, there exists a simple closed form expression for the optimal output weights. As such, training RC networks is significantly faster than other recurrent neural networks such as LSTM networks.
@@
The updates (19) are designed so that the reservoir state $\mathbf{r}(t_{i})$ synchronizes with the input time series of the observations $\mathbf{y}(t_{i})$. As a result, the long-term dynamics of the reservoir should be independent of the initial state $\mathbf{r}(t_{0})$ of the reservoir. This feature is referred to as the echo state property [41, 42]. To ensure that the echo state property is satisfied, we use a commonly used heuristic, namely that the spectral radius of the reservoir matrix $W_{R}$ is less than one [43]. We enforce this property by first defining a sparse matrix $W_{R}$ whose nonzero entries are randomly drawn from the uniform distribution U([0.5,0.5])U([-0.5,0.5]). We subsequently rescale this matrix to ensure the desired spectral radius. Similarly, the entries of the input matrix $W_{in}$ and the biases $\mathbf{b}$ are drawn at random from the uniform distribution U([0.5,0.5])U([-0.5,0.5]). The spectral radius of the input weights $W_{in}$ does not play an important role, and therefore no rescaling is needed for $W_{in}$.
To summarize, the observational time series $\mathbf{y}(t_{i})$ act as an external forcing term that drives the reservoir dynamics (19). The reservoir state $\mathbf{r}(t_{i})$ is mapped to the outputs through the weight matrix $W_{out}$. The outputs are then mapped to the kernel vector $\mathbf{z}(t_{i})$ through the kernel matrix $Z$. Finally, we use this kernel vector in (7) to arrive at the S-DEIM estimate $\widetilde{\mathbf{u}}(\mathbf{z}(t_{i}))$.
Across the examples presented in Section 4, the training process takes under a second. Once the RC network is trained, the S-DEIM estimate is obtained within milliseconds. The precise training and inference times are reported in the bottom half of Table 2. These numbers are obtained by averaging over ten independent runs on a MacBook Pro with Apple M3 processors. The code is written in Matlab version R2023a. The reservoir sparsity refers to the portion of the reservoir matrix $W_{R}$ that is nonzero.